### v1

# GW231123: a product of successive mergers from $\sim 10$ stellar-mass black holes

Abstract
GW231123 is an exceptionally massive binary black hole (BBH) merger with unusually high component spins. Such extreme properties challenge conventional stellar evolution models predicting a black hole mass gap due to pair-instability supernovae. We analyze GW231123 using population-informed priors on BH mass and spin distributions to test possible formation scenarios: first-generation stellar collapse, hierarchical (multi-generation) mergers, and primordial origin. Our analysis strongly prefers scenarios where at least one component is a higher-generation BH. Both components are favored to have high spins, which rules out scenarios in which they are both first-generation (low spin) or primordial (nearly non-spin). We conclude that GW231123 is a hierarchical merger, with components plausibly originate from the successive mergers of $\sim 6$ and $\sim 4$ first-generation BHs, respectively. This suggests that repeated mergers can be frequent and even more massive intermediate-mass black holes may be produced. Thus mechanisms that can efficiently harden the BBHs’ orbits are required, e.g., gas dynamical fraction in the disks of active galactic nucleus.
Binary Black Holes; Gravitational Waves; Stellar Evolution; Active Galactic Nuclei
software: Bilby (Ashton et al., 2019a, version 1.1.4, ascl:1901.011, https://git.ligo.org/lscsoft/bilby/), Dynesty (Speagle, 2020, version 1.0.1, https://github.com/joshspeagle/dynesty), PyMultiNest (Buchner, 2016, version 2.11, ascl:1606.005, https://github.com/JohannesBuchner/PyMultiNest), Precession (Gerosa & Kesden, 2016, version 1.0.3, https://github.com/dgerosa/precession )
{CJK*}
UTF8gbsn
1 Introduction
Very recently, the LIGO-Virgo-KAGRA (LVK) collaboration reported the detection of GW231123 (The LIGO Scientific Collaboration et al., 2025), a binary black hole (BBH) merger with a remarkably large total source-frame mass of $\sim 190$-$265\,M_{\odot}$ and component dimensionless spins of $\chi_{1}\approx 0.9$ and $\chi_{2}\approx 0.8$. Such extreme masses and spins are unprecedented in previous LVK observations (e.g., Abbott et al., 2023a), and they provide a unique opportunity to probe the formation and evolution of intermediate-mass black holes (IMBHs) (Miller & Colbert, 2004; Liang et al., 2017; Abbott et al., 2022).
Theories of stellar evolution suggest that stars with helium core masses up to $\sim 135\,M_{\odot}$ can only leave behind black holes (BHs) lighter than $\sim 65\,M_{\odot}$, due to (pulsational) pair-instability supernovae ((P)PISN; Fowler & Hoyle, 1964; Barkat et al., 1967; Woosley, 2017; Woosley & Heger, 2021). This leads to a so-called pair-instability mass gap (PIMG) in the BH mass spectrum, although the exact edges of the gap remain uncertain currently (Woosley, 2017; Giacobbo et al., 2018; Woosley, 2019; Farmer et al., 2019, 2020; Belczynski et al., 2020). However, a star with a helium core mass $\gtrsim 135\,M_{\odot}$ is expected to directly collapse into an IMBH, providing a possible origin for the components of GW231123.
Alternatively, in dynamical formation channels BHs can undergo repeated mergers and substantially grow in mass, as long as the merger remnants are efficiently retained in their host environments (see Gerosa & Fishbach, 2021; Zevin & Holz, 2022; Li & Fan, 2025a). GW231123 may thus represent a hierarchical merger. However, how massive BHs can grow via hierarchical assembly remains unclear and likely depends on the host environment. For example, in active galactic nucleus (AGN) disk channels the maximum BH mass may be limited by the finite lifetime of the disk (Xue et al., 2025). Additionally, primordial BHs (PBHs) formed in the early Universe (Carr & Hawking, 1974; Khlopov, 2010) could also populate mass ranges that are hard to reach via stellar collapse (Cai et al., 2018; Escrivà et al., 2024).
Previously, the BBH merger GW190521 (Abbott et al., 2020a) attracted significant attention, as it contained at least one component apparently above the lower edge of PIMG. Fishbach & Holz (2020) and Nitz & Capano (2021) proposed that GW190521 might have been a straddling binary, meaning one BH lay below and the other above the PIMG, so that neither BH was born in the disallowed mass range. However, hierarchical merger scenarios were later found to be more natural for GW190521 once spin information was taken into account. In particular, population analyses inferred that at least one component of GW190521 was likely a higher-generation BH based on its spin and mass properties (Wang et al., 2021b; Kimball et al., 2021; Wang et al., 2022; Li et al., 2024b; Antonini et al., 2025b; Li et al., 2025; Antonini et al., 2025a).
The formation history of a GW event is encoded in the observed source parameters. By performing parameter estimation (inferring the component masses, spins, etc.), one can test various astrophysical origin hypotheses. However, the marginal distributions of these parameters often have large uncertainties (Abbott et al., 2019, 2020b, 2021, 2023a), making it challenging to determine the origins of these events from standard analyses alone. Crucially, some source parameters are correlated or degenerate with each other (Pürrer et al., 2016; Tiwari et al., 2018). For example, adopting a particular prior for the spin distribution can influence the inferred mass distribution for the same event. We therefore employ population-informed priors–priors motivated by the astrophysical distribution of BH spins and masses–to better identify the most plausible formation scenario for GW231123. By examining the joint mass-spin posterior under different prior assumptions, we can potentially rule out formation channels that would fail to reproduce the observed joint mass-spin distribution (even if those channels might naively appear consistent with the one-dimensional mass or spin posteriors under default priors).
In this work, we investigate the origins of GW231123’s two components and the formation channel of the system by analyzing its mass and spin properties with population-informed priors. The rest of the paper is organized as follows: In Section 2, we define the origin scenarios and their expected mass-spin signatures, and describe our population-informed prior choices. In Section 3, we present the results of applying these priors to GW231123 and compare the evidence for each scenario. Besides, we also estimate the number of progenitors needed to generate GW231123 in the hierarchical merger scenario. Finally, we discuss the implications and draw conclusions in Section 4.
2 Population Information to Test Formation Hypotheses
In this section, we introduce the population-informed priors used to represent different formation hypotheses for GW231123.
The natal spins of BHs are expected to differ based on their formation mechanism. Stellar-collapse (first-generation) BHs are generally predicted to be born with low spins ($\chi\lesssim 0.1$; Fuller & Ma, 2019), although binary interactions can spin them up modestly (Qin et al., 2018; Bavera et al., 2020; Shao & Li, 2022). In contrast, BHs formed by previous BH-BH mergers tend to have significantly larger spins, typically peaking around $\chi\sim 0.7$ (Gerosa & Berti, 2017; Fishbach et al., 2017; Gerosa & Fishbach, 2021). Indeed, recent population studies of GWTC-3 have found evidence for two distinct subpopulations in the plane of spin magnitude versus component mass (Li et al., 2024b, a; Pierra et al., 2024). The first subpopulation (with smaller BH masses and $\chi\lesssim 0.3$) is consistent with BHs born from stellar core-collapse, whereas the second subpopulation (with more massive BHs and $\chi\sim 0.75$) is indicative of BHs that have undergone at least one merger already.
Motivated by these findings, we introduce spin-magnitude priors corresponding to the low-spin (first-generation) and high-spin (higher-generation) BH populations. Specifically, we take $P_{\rm LS}(\chi)\sim\mathcal{G}(\mu=0.15,\sigma=0.2)$ and $P_{\rm HS}(\chi)\sim\mathcal{G}(\mu=0.8,\sigma=0.3)$, where $\mathcal{G}$ denotes a Gaussian distribution (truncated to the physical range $[0,1]$). These choices are consistent with the spin magnitude distributions inferred from the GWTC-3 BBH population (Li et al., 2024b). We use $P_{\rm LS}$ (“low-spin”) to represent spins of BHs from stellar collapse, and $P_{\rm HS}$ (“high-spin”) to represent spins of BHs formed via previous mergers. For the default prior of spin magnitude, we set $P_{\phi}(\chi)\sim\mathcal{U}(0,1)$, aligning with that of (The LIGO Scientific Collaboration et al., 2025), where $\mathcal{U}$ is the uniform distribution.
Figure 1: Spin magnitude vs. component mass distribution of GW231123 compared to those of BBH events in GWTC-3 informed by population model of Li et al. (2024b). The purple and blue stars (shaded areas) denote the mean values ($90\%$ credible intervals) for the primary and secondary components of GW231123. The green shaded regions show the $90\%$ credible regions for component BHs in the GWTC-3 population, and the black (orange) points mark the mean component masses for the primary (secondary) BHs in those events.
Theoretical predictions also suggest that primordial BHs would have been born with effectively negligible spins (Chiba & Yokoyama, 2017; Green & Kavanagh, 2021). To account for a possible primordial-origin scenario, we include an additional spin prior $P_{\rm NS}(\chi)\sim\mathcal{G}(\mu=0,\sigma=0.01)$, representing a nearly zero-spin distribution. (Whether there exists a subpopulation of BBHs with vanishing spins in current catalog remains under debate (Galaudage et al., 2021; Callister et al., 2022; Tong et al., 2022), but we consider this extreme case for completeness.)
Component masses are also critical to determine the origins of the BHs. Stellar-evolution theory predicts a dearth of BHs in the PIMG roughly between $\sim 40$-$60\,M_{\odot}$ and $\sim 120$-$130\,M_{\odot}$ (e.g., Woosley, 2017; Giacobbo et al., 2018; Woosley, 2019; Farmer et al., 2019, 2020). The precise boundaries of this gap are uncertain and sensitive to model details (nuclear reaction rates, stellar rotation, etc.), but its width is expected to be on the order of $80~{}M_{\odot}$ (Farmer et al., 2019). Some GW population analyses suggest the lower edge of the gap could be as low as $\sim 45~{}M_{\odot}$ (Wang et al., 2021b, 2022; Li et al., 2024b; Pierra et al., 2024; Antonini et al., 2025b). Based on these insights, we impose the following mass priors for stellar-formed BHs: $P_{\rm SM}(m)\sim\mathcal{U}(5,45)M_{\odot}$ and $P_{\rm IM}(m)\sim\mathcal{U}(125,300)M_{\odot}$. We use $P_{\rm SM}$ (“stellar-mass”) and $P_{\rm IM}$ (“intermediate-mass”) for BH masses below and above the PIMG, respectively. In this framework, any BH with mass in the range $\sim 45$ $125~{}M_{\odot}$ cannot be produced by ordinary stellar collapse and is presumed to have a non-stellar origin (most likely a hierarchical merger). Different from that of The LIGO Scientific Collaboration et al. (2025), in our inferences, the priors for component masses are uniform distributions in source frame. For the default prior of component mass, we set $P_{\phi}(m)\sim\mathcal{U}(5,300)$.
We note that a purely non-astrophysical origin (both BHs being primordial) (Carr & Hawking, 1974; Khlopov, 2010; Cai et al., 2018; Escrivà et al., 2024) would in principle allow masses in a much broader range than the stellar paradigm. However, because the PBH mass spectrum is not well constrained (De Luca et al., 2021; Yuan et al., 2024), we do not assign a distinct PBH-specific mass prior. Instead, the PBH scenario is tested primarily through the spin prior $P_{\rm NS}(\chi)$, since PBHs would be characterized by nearly zero spins.
Different formation channels for BBHs predict different spin orientation distributions. Isolated field binary evolution (with processes like mass transfer and tidal locking) (see Mandel & Farmer, 2022, and the reference therein) tends to produce BBHs with nearly aligned spins (i.e., small tilt angles between the BH spins and the orbital angular momentum) (Rodriguez et al., 2016). Especially, the chemically homogeneous evolution (CHE) channel, which is potentially able to produce component BHs as massive as GW231123 (Marchant et al., 2016). In contrast, dynamical formation in dense environments (e.g., multi-body interactions in star clusters) typically leads to isotropic spin orientations (random spin tilt angles) (Talbot & Thrane, 2017). An exception is the dynamical formation in AGN disks, where gas torques can align the spins with the orbital axis (Yang et al., 2019; Tagawa et al., 2020a; Li et al., 2022b). In fact, population studies suggest that a fraction of BBH mergers–especially the more massive, hierarchical ones–may occur in AGN environments (Wang et al., 2021a; Gayathri et al., 2021; Wang et al., 2022; Li & Fan, 2025b; Li et al., 2025; Zhu & Chen, 2025).
However, the spin tilts of GW231123 remain poorly constrained: different waveform models yield inconsistent orientation measurements, and independent analyses of LIGO Livingston and Hanford data also disagree (The LIGO Scientific Collaboration et al., 2025). Therefore, we do not implement a population-informed prior for the spin tilt angles for this event.
3 Results
Table 1: Bayes factors of various population-informed priors (x) compared by the default priors ($P_{0}$)
$\ln\mathcal{B}^{P_{0}}_{x}$ $a_{1}\sim P_{\rm NS}(a)$, $a_{1}\sim P_{\rm LS}(a)$, $a_{1}\sim P_{\rm HS}(a)$, $a_{1}\sim P_{\rm LS}(a)$,
$m_{1}\sim P_{\phi}(m)$ $m_{1}\sim P_{\phi}(m)$ $m_{1}\sim P_{\phi}(m)$ $m_{1}\sim P_{\rm IM}(m)$
$a_{2}\sim P_{\rm NS}(a)$,$m_{2}\sim P_{\phi}(m)$ 26.5 15.1 2.8 17.3
$a_{2}\sim P_{\rm LS}(a)$,$m_{2}\sim P_{\phi}(m)$ 25.4 12.8 1.7 11.1
$a_{2}\sim P_{\rm HS}(a)$,$m_{2}\sim P_{\phi}(m)$ 5.7 1.5 -0.8 1.8
$a_{2}\sim P_{\rm LS}(a)$,$m_{2}\sim P_{\rm IM}(m)$ 25.0 12.0 1.3 11.2
$a_{2}\sim P_{\rm LS}(a)$,$m_{2}\sim P_{\rm SM}(m)$ 25.4 11.9 1.5 12.3
Note: The Bayes factors with $\ln\mathcal{B}^{P_{0}}_{x}<0$ ($\ln\mathcal{B}^{P_{0}}_{x}<2.3$) are marked in blue (orange), denoting that
the $x$ prior is favored (not ruled out), comparing to the default prior $P_{0}$.
3.1 What is the most plausible origin for GW231123?
Table 1 summarizes the natural logarithm of the Bayes factors for all population-informed priors relative to the default priors. We find that a zero-spin prior is disfavored for either component of GW231123, suggesting that a PBH origin (which would predict non-spinning BHs) is unlikely. Furthermore, a low-spin prior is less favored either. In other words, neither component of GW231123 appears to come from the low-spin subpopulation identified by Li et al. (2024b), which is thought to consist of first-generation (stellar-collapse) BHs.
We find the scenarios that the secondary BH belongs to the low-spin, low-mass subpopulation can not be ruled out. However, the primary component should belong to the high-spin subpopulation. This means that while it is possible for the two components of GW231123 to straddle the PIMG similar to the case of GW190521 (Fishbach & Holz, 2020), the primary BH is not formed directly via the core collapse of two massive stars. Additionally, the case the secondary BH being a low-spin IMBH (consistent with stellar-collapse origin) is also possible, as long as the primary BH is highly spinning (consistent with merger remnant).
All of our results indicate that at least one component of GW231123 belongs to the high-spin subpopulation (identified in GWTC-3, Li et al., 2024b), which is consistent with a second- or higher-generation BH origin. This strongly supports the interpretation that GW231123 is a hierarchical merger. We plot the component-mass versus spin-magnitude distributions of GW231123111The posterior samples are adopted from https://zenodo.org/records/16004263, comparing to that of the previous events reweighed [^2] The data and codes can be download from GitHub: Stellar-formed V.S. Merger-formed by the population model in Li et al. (2024b), see Figure 1. It shows that both components are consistent with the high-spin subpopulation, which strongly supports that both components of GW231123 are higher-generation BHs.
Figure 2: Mass distributions of merger remnants formed from first-generation BHs in dynamical formation channels, compared to the component masses of GW231123. The first-generation BH mass distribution is taken from Li et al. (2024a). The pink and grey shaded regions indicate GW231123’s primary and secondary mass distributions, which align with mass distribution of remnants assembled from $\sim 6$ (green) and $\sim 4$ (blue) first-generation BHs, respectively. A self-consistency check comparing these model predictions to the observed higher-generation BH population is provided in Appendix B (see Figure 4).
3.2 How many stars are needed to generate GW231123?
Theoretical predictions and population analysis suggest that isolated binary evolution channels contribute to the LVK’s BBH mergers (see e.g. Zevin et al., 2021; Wang et al., 2022; Godfrey et al., 2023; Li et al., 2024a), which are not expected to participate in subsequent hierarchical mergers (Gerosa & Fishbach, 2021). Only the BHs in dynamical formation channels may undergo hierarchical mergers, and potentially contribute to the GW231123-like events. Recent population analysis revealed a subpopulation of BHs, which dominates the mass range $\sim[20,40]~{}M_{\odot}$ (see e.g. Wang et al., 2022; Ray et al., 2024; Li et al., 2024a), consistent with dynamical formation channels predicted by simulations (see e.g., Antonini & Gieles, 2020; Antonini et al., 2023). Therefore, in estimating how many smaller BHs are required to build GW231123’s components, we adopt the mass distribution of the BH subpopulation associated with first-generation BHs in dynamical formation channels, as identified by Li et al. (2024a) [^3] Data and codes are available at GitHub: Field V.S. Dyanmical. Figure 2 shows the mass distribution of hierarchical merger remnants produced by BHs drawn from this subpopulation, compared against the observed masses of GW231123. From these distributions, we infer that producing a BH of GW231123’s primary mass would require merging roughly 6 first-generation BHs, while the secondary could be built from about 4 first-generation BHs (implying a total equivalent of $\sim 10$ stellar-mass BHs to assemble the system’s $\sim 250~{}M_{\odot}$ total mass).
For consistency, we also compare these remnant distributions with the subpopulation of higher-generation BHs observed in GWTC-3 (Li et al., 2024b, a). We find that the distribution of higher-generation BH masses is consistent with remnants produced by the mergers with $\gtrsim 2$ first-generation BHs, supporting the idea that many of the heavy BHs in GWTC-3 (and GW231123 itself) are products of hierarchical growth (see Figure 4 in Appendix B for a self-consistency check).
The primary component of GW231123 could also arise from a direct merger between a $\sim 125M_{\odot}$ stellar-collapse (1G) IMBH and a $\sim 35M_{\odot}$ BH, which would shorten the merger chain to generate GW231123. However, the predicted final spin in such a scenario is significantly lower than the measured primary spin of GW231123, see Figure 3 (right panel). Consequently, it is more plausible that both components of GW231123 were assembled through successive mergers of first-generation BHs below the PIMG.
Figure 3: Comparison of GW recoil kick velocities (left panel) and final remnant spins (right panel) for mergers with BHs of different generations. The left panel shows the kick velocity distributions, indicating that hierarchical mergers involving higher-generation component BHs produce significantly larger kicks than 1G+1G mergers. The right panel shows the corresponding distributions of the remnants’ final spin magnitudes, with shaded areas comparing the results of GW231123. The 1G+1G, nG+1G, and nG+nG BBHs samples are drawn from the posterior population distribution in Li et al. (2024b).
4 Conclusions and Discussion
GW231123 is the most massive BBH merger detected by LVK to date (Abbott et al., 2023a), making it one of the most intriguing events for tests of astrophysical and fundamental physics (The LIGO Scientific Collaboration et al., 2025).
In this work, we analyzed the data of GW231123 under a variety of population-informed priors corresponding to different formation and evolutionary scenarios. Our results show that neither component of GW231123 is consistent with being non-spinning, effectively ruling out a PBH origin (since primordial BHs are expected to have essentially zero spin in certain models; e.g., Chiba & Yokoyama 2017). A low-spin origin for the GW231123 components is also disfavored–especially for the primary BH. Notably, the scenario in which the secondary BH belongs to the low-spin, low-mass subpopulation (i.e. is a first-generation BH, as defined in Li et al. 2024b) cannot be entirely excluded if the primary BH is highly spinning. This suggests that GW231123 could be a “straddling” binary straddling the PIMG, although it likely did not form from the direct collapse of two massive stars (as was once proposed for GW190521; Fishbach & Holz 2020).
Overall, both components of GW231123 are best described by the high-spin subpopulation, consistent with the higher-generation BHs (Gerosa & Fishbach, 2021; Li et al., 2024b). Moreover, the component masses of GW231123 are exceptionally large compared to those in previous BBH observations–particularly in the case of the primary. In fact, the primary’s mass (and might the secondary’s as well) is more than twice the commonly assumed lower bound of the PIMG. Besides, the spin magnitudes of the two components tend to be more consistent with the remnants of mergers with higher-generation BHs, as shown in Figure 3 (right panel). Therefore, it is more likely that both components of GW231123 have undergone successive mergers with several first-generation (stellar-collapse) BHs.
Using the distribution of first-generation BH masses from dynamical formation channels identified in GWTC-3 by Li et al. (2024a) (see also Wang et al. 2022; Ray et al. 2024), we estimate that the primary (secondary) BH in GW231123 could be built up by the merger of $\sim 6$ ($\sim 4$) first-generation BHs. The assembly of such a massive system via hierarchical mergers requires that two key conditions be met. First, earlier merger remnants must be retained in the same environment to undergo subsequent mergers; this means the host’s escape velocity must exceed the gravitational recoil (kick) velocity of each merger remnant. Hierarchical BH mergers impart substantially larger kicks than first-generation mergers (see the left panel of Figure 3), so only environments with sufficiently deep potential wells–such as nuclear star clusters or AGN disks–could keep the GW231123 progenitors bound after each merger. Second, the binary orbits must be harden efficiently so that multiple mergers can occur within a reasonable timescale. In dense stellar systems, dynamical interactions (e.g., binary-single encounters) help harden binaries, but an even more efficient mechanism is likely needed for rapid, repeated mergers. Gas dynamical friction in AGN disks can rapidly shrink binary orbits and shorten merger timescales (Yang et al., 2019), and simulations suggest that this gas-assisted hardening may dominate the upper limit on BBH merger masses (Vaccaro et al., 2024; Xue et al., 2025). Therefore, among known environments, AGN disks are the more plausible sites for producing BBH mergers as massive as GW231123.
If AGN-disk channels indeed dominate the formation of GW231123-like events, then BBH merger masses up to $\sim\mathcal{O}(10^{4})M_{\odot}$ might be achievable (Vaccaro et al., 2024), which are well beyond the detection range of the current generation of ground-based detectors (LIGO Scientific Collaboration et al., 2015; Acernese et al., 2015; Akutsu et al., 2018). The next generation of gravitational-wave observatories, including planned space-based detectors (Luo et al., 2016; Hu & Wu, 2017; Amaro-Seoane et al., 2017; Abbott et al., 2017; Punturo et al., 2010), will be required to observe such extremely massive mergers, offering a more complete view of the BBH mass spectrum and evolutionary pathways. Beyond standard astrophysical channels, a few exotic formation scenarios have been proposed. For example, “dark stars” (hypothetical early-universe stars powered by dark matter annihilation) could form IMBHs; however, the expected masses of such remnants are far larger than those of GW231123 (Lei et al., 2025a). Another proposal involves a cosmological coupling mechanism that gradually increases BH masses, which in principle might produce BHs as massive as GW231123 (Croker et al., 2021). This cosmological coupling hypothesis, however, has been ruled out by very recent JWST observations (Lei et al., 2025b). Ultimately, a comprehensive population analysis of BBH mergers will be necessary to fully elucidate the origins and formation channels of GW231123 and similar exceptional events (e.g., Li et al., 2021, 2022a; Abbott et al., 2023b; Li et al., 2024b; Guo et al., 2024; Alvarez-Lopez et al., 2025).
Acknowledgments
This work is supported by the National Natural Science Foundation of China (No. 12233011, No. 12203101, and No. 12303056), the General Fund (No. 2024M753495) of the China Postdoctoral Science Foundation, and the Priority Research Program of the Chinese Academy of Sciences (No. XDB0550400). This research has made use of data and software obtained from the Gravitational Wave Open Science Center (https://www.gw-openscience.org), a service of LIGO Laboratory, the LIGO Scientific Collaboration and the Virgo Collaboration. LIGO is funded by the U.S. National Science Foundation. Virgo is funded by the French Centre National de Recherche Scientifique (CNRS), the Italian Istituto Nazionale della Fisica Nucleare (INFN) and the Dutch Nikhef, with contributions by Polish and Hungarian institutes.
Appendix A Bayesian Inference
We analyze an 8-s segment of strain data for GW231123, spanning from $t_{\rm c}-6{\rm s}$ to $t_{\rm c}+2{\rm s}$ (where $t_{\rm c}$ is the geocentric GPS time of coalescence), which are available from the Gravitational Wave Open Science Center Catalog [^4] Download from https://doi.org/10.5281/zenodo.15832843. For the noise power spectral density (PSD), we use the pre-estimated PSD files released by The LIGO Scientific Collaboration et al. (2025).
Our primary waveform model is the precessing multipolar approximant IMRPhenomXPHM (Khan et al., 2019; Hannam et al., 2014; Pratten et al., 2021), which includes subdominant harmonics. As a cross‐check, we also perform inference with IMRPhenomXO4a (Thompson et al., 2024), which was shown to exhibit a distinct secondary mode in the component‐mass posterior in The LIGO Scientific Collaboration et al. (2025). The corresponding results for this waveform model are presented in Table 2.
[…]
where we take $f_{\rm min}=20$ Hz and $f_{\rm max}=488$ Hz following The LIGO Scientific Collaboration et al. (2025). Except for the population-informed priors introduce in Section 2, all other priors follow those of The LIGO Scientific Collaboration et al. (2025). We sample the posterior distributions using the Bilby (Ashton et al., 2019b) framework with PyMultiNest (Buchner, 2016) sampler, and verify results with Dynesty (Speagle, 2020).
Table 2: Bayes factors obtained with IMRPhenomXO4a
$\ln\mathcal{B}^{P_{0}}_{x}$ $a_{1}\sim P_{\rm NS}(a)$, $a_{1}\sim P_{\rm LS}(a)$, $a_{1}\sim P_{\rm HS}(a)$, $a_{1}\sim P_{\rm LS}(a)$,
$m_{1}\sim P_{\phi}(m)$ $m_{1}\sim P_{\phi}(m)$ $m_{1}\sim P_{\phi}(m)$ $m_{1}\sim P_{\rm IM}(m)$
$a_{2}\sim P_{\rm NS}(a)$,$m_{2}\sim P_{\phi}(m)$ 40.1 11.6 -0.5 11.1
$a_{2}\sim P_{\rm LS}(a)$,$m_{2}\sim P_{\phi}(m)$ 22.9 10.2 0.3 10.3
$a_{2}\sim P_{\rm HS}(a)$,$m_{2}\sim P_{\phi}(m)$ 8.0 8.3 -1.2 7.5
$a_{2}\sim P_{\rm LS}(a)$,$m_{2}\sim P_{\rm IM}(m)$ 23.1 9.0 -2.2 12.4
$a_{2}\sim P_{\rm LS}(a)$,$m_{2}\sim P_{\rm SM}(m)$ 19.9 9.3 -3.2 9.9
Note: The Bayes factors with $\ln\mathcal{B}^{P_{0}}_{x}<0$ ($\ln\mathcal{B}^{P_{0}}_{x}<2.3$) are marked in blue (orange), denoting that
the $x$ prior is favored (not ruled out), comparing to the default prior $P_{0}$.
Appendix B Self-consistency check for BH generations
To validate our choice of the first-generation mass distribution in dynamical formation channels, we compare the mass distribution of remnants made by multiple ($n$) first-generation BHs with those of subpopulations revealed in Li et al. (2024a) and Li et al. (2024b). As shown in the top panel of Figure 4, if we assume the first-generation BHs are drawn from the potential dynamical 1G subpopulation (green shaded area) revealed in Li et al. (2024a), remnants from $\gtrsim 2$ progenitors can well reproduce the higher-generation subpopulation (blue shaded area). Additionally, the higher-generation BHs are dominated by the remnants made by 2 first-generation BHs, which is natural for the scenario in hierarchical mergers. However, if we assume the first-generation BHs are drawn from the low-spin subpopulation (orange shaded area) as revealed in Li et al. (2024b), then the mass distribution of remnants made by 2 first-generation BHs would peak at $\sim 20M_{\odot}$, which is inconsistent with the distribution of the higher-generation subpopulation (blue shaded area), see the bottom panel of Figure 4. This discrepancy further supports our adoption of the dynamical first-generation subpopulation for modeling hierarchical growth.
In practice, we do not account for the mass loss due to GW radiation. Because GW emission only reduces the remnant mass by a few percent, an effect small compared to current measurement uncertainties. We also neglect mass accretion in plausible gas‐rich environments (e.g., AGN disks), which could increase BH masses by $\sim 10\%$-$20\%$ (Tagawa et al., 2020b; Xue et al., 2025; Vaccaro et al., 2024). Including these effects would not alter our qualitative conclusions but would slightly change the number (at most one) of first‐generation progenitors required to assemble GW231123.
Figure 4: Mass distributions of the subpopulations comparing to the remnants of mergers drawn from the potential first-generation subpopulation. Top: the first-generation BHs are drawn from the potential dynamical first-generation subpopulation (green region), the subpopulation in purple are potentially associated with field evolution channels. Bottom: the first-generation BHs are drawn from the low-spin subpopulation (orange region), which is likely the mixture of dynamical and field channels. Note that each distribution of subpopulation is normalized, and the shaded regions are for $90\%$ credible levels.